OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph.

OpenStudy (anonymous):

To find the slope, we ned to put it in slope intercept form: y=-x+5 The slope is -1

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

hence the slope(m)=1

OpenStudy (anonymous):

Given the linear equation 3x + y = 5, find the slope of its graph.

OpenStudy (anonymous):

m=3

OpenStudy (anonymous):

hence the slope is 3

OpenStudy (anonymous):

y=-3x+5 slope is -3

OpenStudy (anonymous):

y=-3x+5

OpenStudy (anonymous):

Given the linear equation 2y + 6 = 5x, find the y-intercept.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Geez, do it yourself. A lot fo ppl are helping u and still u dont get it

OpenStudy (anonymous):

do u understand how to do it?????

OpenStudy (anonymous):

be nice @powerful22

OpenStudy (anonymous):

im checking my answers!!

OpenStudy (anonymous):

2y=5x-6 y=5/2x-3 slope is 5/2

OpenStudy (anonymous):

ok, good!

OpenStudy (anonymous):

Given the linear equation 3y + 4x =12, find the y-intercept.

OpenStudy (anonymous):

at least show you're answer so we can verify it

OpenStudy (anonymous):

yes the slope is 5/2 very good @zeinas.

OpenStudy (anonymous):

there for my study guide for my tests next week

OpenStudy (anonymous):

thanks :) @_CALCULASMASTER_

OpenStudy (anonymous):

Given the linear equation 3y + 4x =12, find the y-intercept.

OpenStudy (anonymous):

yah np

OpenStudy (anonymous):

can u help with this last one??

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

3y+4x=12 3y=-4x+12 y=-4/3x+4 y-intercept is 4

OpenStudy (anonymous):

Given the linear equation 3y + 4x =12, find the y-intercept.

OpenStudy (anonymous):

y=mx+b 3y + 4x = 12 3y =-4x + 12 y = -4/3x + 4 y=mx+b m=slope b=y-intercept so what is the answer?